# Danny Halawi4, Fred Zhang4, Philip E. Tetlock2,5

### v4

@@ Appendix H Leaderboards: Top 50
5. Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
Appendix I Human Survey Demographics
@@
We collected demographic information from the 500 human forecasters in the public survey. Summaries of participants’ age, gender, ethnicity, and country of residence are shown in the tables below.
@@
18-24 years old 32.0%
25-34 years old 43.4%
35-44 years old 14.4%
45-54 years old 5.4%
@@
We did not collect similar demographic information from the superforecasters participating in the study, but are reasonably certain that the superforecasters in this study are roughly representative of superforecasters as a whole. Describing forecasters previously recruited by Good Judgment Project, Mellers et al. (2015) noted that participants "tended to be men (83%) and U.S. citizens (74%), with an average age of 40 years."
Appendix J Performance Breakdown
In tables Table 25 and Table 26 we look into how the best-performing LLM, Claude 3.5 Sonnet (Scratchpad prompt with freeze values), and Superforecasters performed when assessing performance by category and by forecast horizon.
In Table 25, we see that Claude 3.5 Sonnet slightly Superforecasters on Environment & Energy questions and Sports questions.
As a reminder, we ask for forecasts on dataset questions at 8 diffreent forecast horizons, as explained in Section 3.2. Here, we breakdown performance at the forecast horizons that have resolved by the publication date: 7 days, 30 days, and 90 days. We see that Superforecasters outperformed Claude 3.5 Sonnet at all forecast horizons.
@@
Arts & Recreation 8 0.086 0.053 0.033
Economics & Business 137 0.188 0.105 0.083
Environment & Energy 63 0.145 0.158 -0.013
Healthcare & Biology 37 0.003 0.001 0.002
Politics & Governance 17 0.061 0.017 0.044
Science & Tech 15 0.160 0.139 0.021
Security & Defense 73 0.049 0.027 0.022
Sports 43 0.020 0.034 -0.014

### v5

@@ Appendix H Leaderboards: top 50
5. Pairwise $p$-value comparing to No. 1 (bootstrapped): The $p$-value calculated by bootstrapping the differences in overall score between each model and the best
forecaster under the null hypothesis that there’s no difference.
6. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster.
Appendix I Aggregating human forecasts
We provide aggregated forecasts on benchmark questions from 500 members of the general public and 39 superforecasters, as described in Section 4:
As described in Section 4, members of the general public were recruited via Prolific and Facebook. First, participants completed an introductory survey designed to gather demographic information and evaluate performance on a few forecasting and comprehension tasks. Then, they were invited to take part in the main survey, featuring 20 random benchmark questions. Some participants were disqualified from participating in the main survey based on suspicious elements of their presurvey responses (e.g., answering questions rapidly, large numbers of submissions from the same IP address, etc.).
Superforecasters were invited to participate directly and were prompted to give forecasts for at least 20 questions from the benchmark (though, many chose to participate on additional questions). Superforecasters were also given access to other Superforecasters’ forecasts and rationales and were allowed to comment on and update based on others’ forecasts.
For each group, the median forecast for each forecasting question was taken to create the aggregated forecast sets. Code and individual forecasts are forthcoming once we can ensure that the text responses have been fully anonymized.
Appendix J Reproduce LLM forecasts
We evaluate 17 LLMs on our initial benchmark: GPT-3.5-Turbo-Instruct [Brown et al., 2020], GPT-4 [OpenAI, 2023], GPT-4o, Llama-2-70B [Touvron et al., 2023], Llama-3-7B, Llama-3-70B, Mistral-7B, Mistral-8x7B [Jiang et al., 2024a], Mistral-8x22B, Mistral-Large, Qwen1.5-110B-Chat, Claude-2.1 [Anthropic, 2023], Claude-3-Haiku, Claude-3.5-Sonnet, Claude-3-Opus [Anthropic, 2024], Gemini 1.5 Flash, and Gemini 1.5 Pro [Gemini Team, 2023].
To make inferences, we use APIs. For the GPT-suite, we use OpenAI’s API; for Gemini-suite, we use Google’s API; for Llama-suite, Mistral-7B, Mixtral-8x7B, Mixtral-8x22B and Qwen1.5-110B-Chat, we use Together AI’s API; for Mistral-Large, we use Mistral AI’s API; and for the Claude-suite, we use Anthropic’s API. To reproduce our results, people will need to gather these API keys.
We then record model predictions using several different methods: zero-shot prompting, scratchpad prompting, scratchpad prompting with retrieval augmentation, and scratchpad prompting with retrieval augmentation and aggregate human forecasts. For the scratchpad and information retrieval setting, we use the retrieval infrastructure from Halawi et al. [2024] and provide relevant news articles to the models in-context to reason about. Additionally, only models with a context window larger than 8,000 tokens were evaluated under the retrieval setting due to the inclusion of news articles in the prompt.
Additionally, to speed up the inference, we use multithreading with 50 workers, which requires a high rate limit and requests for better subscription plans from each source. However, one can run inference sequentially by setting it as 1 worker, but this requires longer time to generate all the baselines.
J.1 Zero-shot and scratchpad baselines
Prompts We use the zero-shot and scratchpad prompts shown in Appendix F.
Hyperparameters For the zero-shot setting, we set the maximum output token length to 50 since we only request probabilistic forecasts. For the scratchpad prompt, we increase the maximum output token length to 1300 as it requires reasoning and probabilistic forecasts. We initially considered a high token length of 3000, but after observing that the maximum response length was around 1250, we settled on 1300 as the optimal maximum token length. In both cases, the model temperature is set to 0 to ensure stable outputs.
How to Reproduce To run zero-shot and scratchpad baselines, follow the steps below:
  1. 1.
    Insert all the necessary API keys in src/helpers/keys.py.
  2. 2.
    Run the zero-shot and the scratchpad baselines in src/base_eval/llm_baselines/
    manager/main.py
J.2 Scratchpad with information retrieval baseline
Prompts We use the same scratchpad prompt as scratchpad baseline with an additional line “We have retrieved the following information for this question: {retrieved_info}” before the instructions begin.
Information Retrieval We use the same information retrieval system from Halawi et al. [2024]. The pipeline consists of four steps: search query generation, news retrieval, relevance filtering and re-ranking, and text summarization. One must acquire a Newscatcher API key to implement the same retrieval method.
Information Retrieval Hyperparameters The hyperparameters were selected following the results in Section E.1 of Halawi et al. [2024], in which they used a greedy search approach to identify the optimal hyperparameters. We display the hyperparameters below:
NUM_SEARCH_QUERY_KEYWORDS:
The number of keywords used in the search query. For our system, this is set to 6.
MAX_WORDS_NEWSCATCHER:
The maximum number of words allowed in search queries for the NewsCatcher API. This is set to 5.
MAX_WORDS_GNEWS:
The maximum number of words allowed in search queries for the Google News API. This is set to 8.
SEARCH_QUERY_MODEL_NAME:
The name of the model used to generate search queries. We use gpt-4-1106-preview.
SEARCH_QUERY_TEMPERATURE:
The temperature setting for the search query model, which controls the randomness of the output. We set this to 0.0 for deterministic outputs.
SEARCH_QUERY_PROMPT_TEMPLATES:
The templates used to generate search queries. In our configuration, we use PROMPT_DICT["search_query"]["0"] and PROMPT_DICT["search_query"]["1"]. The exact search query can be found in search_query.py.
NUM_ARTICLES_PER_QUERY:
The number of articles retrieved per search query. This is set to 10.
SUMMARIZATION_MODEL_NAME:
The name of the model used for summarizing articles. We use gpt-3.5-turbo-1106.
SUMMARIZATION_TEMPERATURE:
The temperature setting for the summarization model, which controls the randomness of the output. We set this to 0.2.
SUMMARIZATION_PROMPT_TEMPLATE:
The template used for summarizing articles. In our configuration, we use PROMPT_DICT["summarization"]["9"]. The exact search query can be found in summarization.py.
NUM_SUMMARIES_THRESHOLD:
The threshold number of summaries to generate. This is set to 10.
PRE_FILTER_WITH_EMBEDDING:
A boolean flag indicating whether to pre-filter articles using embeddings. This is set to True.
PRE_FILTER_WITH_EMBEDDING_THRESHOLD:
The threshold for pre-filtering articles using embeddings. This is set to 0.32.
RANKING_MODEL_NAME:
The name of the model used for ranking articles. We use gpt-3.5-turbo-1106.
RANKING_TEMPERATURE:
The temperature setting for the ranking model, which controls the randomness of the output. We set this to 0.0 for deterministic outputs.
RANKING_PROMPT_TEMPLATE:
The template used for ranking articles. In our configuration, we use PROMPT_DICT["ranking"]["0"].
RANKING_RELEVANCE_THRESHOLD:
The relevance threshold for ranking articles. This is set to 4.
RANKING_COSINE_SIMILARITY_THRESHOLD:
The cosine similarity threshold used in ranking. This is set to 0.5.
SORT_BY:
The criterion used to sort articles. We sort by date.
RANKING_METHOD:
The method used for ranking articles. We use llm-rating.
RANKING_METHOD_LLM:
The specific method for ranking articles using the LLM. We use title_250_tokens, meaning ranking articles based on their titles and the first 250 tokens.
NUM_SUMMARIES_THRESHOLD:
The threshold number of summaries to generate for final output. This is set to 20.
EXTRACT_BACKGROUND_URLS:
A boolean flag indicating whether to extract background URLs from the articles. This is set to True.
Inference Hyperparameters: We set the maximum output token length to 2000 to accommodate reasoning and probabilistic forecasts. We set the model temperature to 0 to ensure stable outputs.
How to reproduce To run the Scratchpad with Information Retrieval baseline, follow these steps:
  1. 1.
    To run the information retrieval part:
    1. (a)
      Insert all the necessary API keys in llm_retrieval/
      forecasting-llm-retrieval/config/keys.py
      . Specifically, add the Newscatcher and OpenAI API keys.
    2. (b)
      Run llm_retrieval/notebooks/retrieval_cache.ipynb.
    3. (c)
      Save all the retrieved news under a folder called news.
  2. 2.
    To run the scratchpad with the information retrieval baseline:
    1. (a)
      Insert all the necessary API keys in src/helpers/constants.py.
    2. (b)
      Place the "news" folder in the same directory as src/base_eval/all_recurrent_llm_baselines/main.py.
    3. (c)
      Run src/base_eval/all_recurrent_llm_baselines/main.py.
J.3 LLM "ensemble" baseline
To produce the LLM ensemble forecast, we query three models: GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Pro. We use three prompts crafted by superforecasters who were given explicit instructions to write prompts that would help an LLM produce accurate forecasts This results in $3\times 3=9$ forecasts per question. We then show 3 LLM crowd baselines using the median, geometric mean, and geometric mean of log odds.
Prompts We use the 3 superforecaster-written prompts shown in the appendix of our paper as Superforecaster Prompt 1-3.
Inference hyperparameters We set the maximum output token length to 2000 to accommodate reasoning and probabilistic forecasts. We initially considered a high token length of 3000, but after observing that the maximum response length was around 1950, we finalized 2000 as the optimal maximum token length. We set the model temperature to 0 to ensure stable outputs.
How to reproduce To run LLM "Ensemble" baseline, follow the below steps:
  1. 1.
    Insert all the necessary API keys in src/helpers/constants.py.
  2. 2.
    Place the "news" folder in the same directory as src/base_eval/llm_crowd/
    notebook.ipynb
  3. 3.
    Run src/base_eval/llm_crowd/notebook.ipynb.
Appendix K Reproduce resolution and leaderboard
Given the forecast files output from Appendix I and Appendix J, the forecasts can be resolved and the leaderboard created as outlined below, after first having downloaded the benchmark codebase.
The Google Cloud Run Job in src/resolve_forecasts/main.py resolves all forecasts on the questions from the question set in Section B.1. To do this, it depends on:
  • the forecast files provided in Section B.2.1 and Section B.2.3;
  • the complete resolution files from our Question Bank on GCP Cloud Storage, which we cannot distribute freely because some providers do not allow us to distribute their data directly, rather only modifications of their data. However, the code to create these resolution files is provided under src/questions and can be created given the API keys to the data sources.
Having resolved the forecasts for the day, either to ground truth if it was a forecast on a dataset question, or the resolution value or market value for market questions, we can now create the leaderboard. To do this, we use the Google Cloud Run Job defined in src/leaderboard/main.py.
Appendix L General public survey demographics
@@ Appendix I Aggregating human forecasts
We collected demographic information from the 500 human forecasters in the general public survey. Summaries of participants’ age, gender, ethnicity, and country of residence are shown in the tables below.
@@
18–24 years old 32.0%
25–34 years old 43.4%
35–44 years old 14.4%
45–54 years old 5.4%
@@
We did not collect similar demographic information from the superforecasters participating in the study, but are reasonably certain that the superforecasters in this study are roughly representative of superforecasters as a whole. Describing forecasters previously recruited by Good Judgment Project, Mellers et al. [2015] noted that they “tended to be men (83%) and U.S. citizens (74%), with an average age of 40 years.”
Appendix M Performance breakdown
Tables Table 25 and Table 26 show the performance of the top LLM—Claude 3.5 Sonnet (using the Scratchpad prompt with freeze values)—compared with Superforecasters, evaluated by forecast category and horizon.
In Table 25, we see that Claude 3.5 Sonnet slightly outperformed superforecasters on Environment & Energy questions.
As a reminder, we ask for forecasts on dataset questions at 8 diffreent forecast horizons, as explained in Section 3.2. Here, we breakdown performance at the forecast horizons that have resolved by the publication date: 7 days, 30 days, 90 days, and 180 days. We see that superforecasters outperformed Claude 3.5 Sonnet at all forecast horizons, except for the 90-day horizon where Claude 3.5 Sonnet performed better.
@@
Arts & Recreation 8 0.102 0.066 0.036
Economics & Business 181 0.224 0.134 0.090
Environment & Energy 81 0.126 0.141 -0.015
Healthcare & Biology 48 0.003 0.001 0.002
Politics & Governance 17 0.066 0.022 0.044
Science & Tech 14 0.207 0.175 0.032
Security & Defense 95 0.056 0.030 0.026
Sports 54 0.058 0.034 0.024